# UDA4Inst: Unsupervised Domain Adaptation for Instance Segmentation*

### v6

@@ Abstract
Instance segmentation is crucial for autonomous driving, but is hindered by the lack of annotated real-world data due to expensive labeling costs. Unsupervised Domain Adaptation (UDA) offers a solution by transferring knowledge from labeled synthetic data to unlabeled real-world data. While UDA methods for synthetic to real-world domains (synth-to-real) excel in tasks such as semantic segmentation and object detection, their application to instance segmentation for autonomous driving remains underexplored and often relies on suboptimal baselines. We introduce UDA4Inst, a powerful framework for synth-to-real UDA in instance segmentation. Our framework enhances instance segmentation through Semantic Category Training and Bidirectional Mixing Training. Semantic Category Training groups semantically related classes for separate training, improving pseudo-label quality and segmentation accuracy. Bidirectional Mixing Training combines instance-wise and patch-wise data mixing, creating coherent composites that enhance generalization across domains. Extensive experiments show UDA4Inst sets a new state-of-the-art on the SYNTHIA → Cityscapes benchmark (mAP 31.3) and introduces results on novel datasets, using UrbanSyn and Synscapes as sources and Cityscapes and KITTI360 as targets. Code and models are available at https://github.com/gyc-code/UDA4Inst.
@@ I Introduction
Furthermore, we apply a bidirectional domain mixing training strategy to provide accurate pseudo-labels from the selected target domain. Unlike the previous model [16] that limited the cross-domain data mixing to a single direction, either source-to-target (S2T) or target-to-source (T2S), UDA4Inst combines and balances both the strategies accurately to address the instance segmentation and provide the pseudo-labels.
Note that one of the most challenging scenarios in instance segmentation is when multiple instances of the same class overlap. In semantic segmentation, these overlapping instances are treated as a single blob. In contrast, instance segmentation methods need to find clear boundaries in the overlapping region of objects to separate them. We solve this challenge by adopting an instance-wise mixing strategy for large-area instances, ensuring that composite images retain clear object boundaries. For smaller instances, we use a patch-wise mixing method, ensuring that the instances in the mixed images retain the surrounding environmental context. This prevents small, indistinct objects from appearing abruptly or unnaturally in the new domain images. Additionally, to address instance imbalance, UDA4Inst incorporates a straightforward online algorithm that balances rare-class instances during data mixing between domains.
@@
We demonstrate the effectiveness of UDA4Inst on multiple UDA instance segmentation benchmarks. On SYNTHIA [17]Cityscapes, we achieve state-of-the-art results. To the best of our knowledge, this is the first report performing UDA instance segmentation using the synthetic driving scene datasets UrbanSyn [18] and Synscapes [19], achieving 38.6 mAP and 35.9 mAP on Cityscapes [20] and 25.0 mAP and 19.5 mAP on KITTI360 [21]. Compared to the source-only method, our approach demonstrates substantial improvements across these benchmarks as in Fig. 1. Our contributions are as follows:
@@ III-A Overall Approach
Our paper introduces the UDA4Inst framework. Fig. 2 provides an overview of the pipeline. We perform Semantic Category Training by dividing the training data into two semantic groups, (e.g. Human-Cycle and Vehicle). We train an individual model for each group. Moreover, we apply the bidirectional mixing training, which involves cross-domain cut-and-paste operations performed both instance-wise and patch-wise to train the model. The outputs of these models are subsequently fused and refined to generate high-quality pseudo-labels from the unlabeled target data. Our final step uses only these pseudo-labels to train a single final model and then produce the final predictions in the target domain. This framework aims to generate high-quality pseudo-labels, which are used to train with the target domain data in a supervised manner, ultimately producing the final instance segmentation results. After an overall explanation of our UDA4Inst framework, we explain key proposals in more detail: Semantic Category Training and Bidirectional Mixing Training.
@@ III-C Bidirectional Mixing Training
On each of the two categories, we apply our Bidirectional Mixing Training. Where the source domain dataset , comprising images , their instance labels , an instance segmentation model ; is trained in a supervised manner. The target domain dataset only provides images . We aim to address the domain gap by leveraging , , and to train a model capable of accurately segmenting instances on .
As shown in Fig. 3, our teacher-student UDA method is based on online self-training. The training has two stages. In the first stage, the model with parameters is trained on the source domain in a supervised manner for iterations, with synthetic images and their corresponding ground-truth labels , to generate predictions . Our goal is to minimize a supervised segmentation loss
@@ III-C2 Overlapping Strategy
Overlapping instances in the mixing process can create ambiguous labels and unrealistic visual artifacts, potentially confusing the model during training. To address this, we implement a priority masking strategy. Specifically, in S2T mixing, we assign higher priority to the source domain instances. If an instance from the source image overlaps with an instance in the target image, we erase the overlapping region of the target’s instance labels, keeping only the source instance in that region. This means that the source instance’s mask replaces any conflicting target instance masks in the overlapping area. We do a similar strategy for T2S mixing.
@@ IV-B1 Training
We adopt Mask2Former [15] with Swin Transformer backbone [37] as our base network . The model is pre-trained on COCO [38] dataset. We apply the same data augmentation (flipping, random cropping, etc.) as [15] and set the image crop size to when the target domain is Cityscapes and for KITTI360. The maximum number of samples N for the rare-class pool is set to 10, and the pseudo-label confidence threshold is 0.9. We set the numbers of iterations for source domain training and mixed data training to be 40k each. During validation, we follow the same checkpoint selection protocol used in [28]. For each stage of training, we set a batch size =3, considering optimal memory usage. We use AdamW [39] optimizer with an initial learning rate of 0.0001 and a weight decay of 0.05, as well as the ‘poly’ learning rate schedule with the power of 1. The hyperparameters of the loss , , and are set to 5.0, 5.0, 2.0 respectively. During the supervised training phase using pseudo-labels, we adopt the same training configuration. For bidirectional mixing, we set the pixel area threshold to 1500; instances with an area larger than this threshold are processed instance-wise, while those below are handled patch-wise. For these experiments, UDA4Inst runs on 1 NVIDIA A40 GPU with 48GB of memory.
@@ IV-C1 UDA performance
In Table I, we show how UDA4Inst improves over the vanilla Mask2Former baseline in synth-to-real settings. We also include the supervised Mask2Former (trained on Cityscapes and KITTI360) as upper bounds. UDA4Inst consistently outperforms the source-only model, achieving mAP of 38.6 (UrbanSynCityscapes), 35.9 (SynscapesCityscapes), and a new state-of-the-art 31.3 (SYNTHIACityscapes)—representing +7.0, +7.8, and +6.7 mAP gains over Mask2Former. Improvements are particularly notable for larger objects like Car, Truck, Bus, and Train, while smaller, more deformable classes (Person, M.bike, and Bike) see modest gains. Fig. 6 further illustrates the advantage of UDA4Inst over the source-only model in UrbanSynCityscapes. UDA4Inst also shows strong performance on KITTI360, reaching 23.2 mAP (UrbanSynKITTI360) and 20.1 mAP (SynscapesKITTI360). Among the source datasets, UrbanSyn consistently performs best on both Cityscapes and KITTI360, likely due to its high photorealism, which narrows the domain gap and boosts adaptation.
@@ IV-D1 Component Ablation
Table III shows how UDA4Inst’s three components—Semantic Category Training (SCT), Bidirectional Mixing Training (BMT), and Rare-class Balancing (RcB)—affect UrbanSynCityscapes. M1 (no UDA) obtains the lowest mAP of 32.0, while M6 (all components) reaches 39.0. Comparing M1 vs. M2 and M3 vs. M4 confirms SCT’s effectiveness, while BMT alone boosts mAP from 32.0 (M1) to 36.1 (M3). With SCT, BMT further improves M2 by +2.9 mAP. Adding RcB (M5) yields substantial gains for rare classes like Train and Motorcycle. Among all components, BMT stands out as the most impactful for adaptation.
@@
w/o mixing | 31.95 |
S2T mixing | 36.07 |
T2S mixing | 35.00 |
bidirectional mixing | 37.18 |
@@ IV-D2 Bidirectional Mixing Training
We examine data mixing strategies without RcB or SCT in Table IV (UrbanSynCityscapes). Both S2T and T2S mixing help, although T2S is slightly weaker due to noise in target pseudo-labels. Combining S2T and T2S achieves 37.18 mAP, a +5.23 gain over the baseline without mixing, balancing target learning with pseudo-label noise mitigation. We also evaluate instance-wise vs. patch-wise mixing in Table V.
@@ IV-D4 Semantic Category Training
Table VII shows that the category approach consistently surpasses the full-class method across all domains. For instance, it yields a +6.3 mAP boost on UrbanSyn and +5.9 on SYNTHIA. Although improvements on Cityscapes are smaller (up to +2.9 mAP), the category model still generalizes better. Similarly, on KITTI360, the category model outperforms the full-class variant. These results confirm that training specialized models for semantically similar classes enhances both source-domain performance and domain adaptation, making semantic grouping a promising strategy for UDA instance segmentation.

### v7

@@ Abstract
Instance segmentation is crucial for autonomous driving, but is hindered by the lack of annotated real-world data due to expensive labeling costs. Unsupervised Domain Adaptation (UDA) offers a solution by transferring knowledge from labeled synthetic data to unlabeled real-world data. While UDA methods for synthetic to real-world domains (synth-to-real) excel in tasks such as semantic segmentation and object detection, their application to instance segmentation for autonomous driving remains underexplored and often relies on suboptimal baselines. We introduce UDA4Inst, a powerful framework for synth-to-real UDA in instance segmentation. Our framework enhances instance segmentation through Semantic Category Training and Bidirectional Mixing Training. Semantic Category Training groups semantically related classes for separate training, improving pseudo-label quality and segmentation accuracy. Bidirectional Mixing Training combines instance-wise and patch-wise data mixing, creating coherent composites that enhance generalization across domains. Extensive experiments show UDA4Inst sets a new state-of-the-art on the SYNTHIA Cityscapes benchmark (mAP 31.3) and introduces results on novel datasets, using UrbanSyn and Synscapes as sources and Cityscapes and KITTI360 as targets. Code and models are available at https://github.com/gyc-code/UDA4Inst.
@@ I Introduction
Furthermore, we apply a bidirectional domain mixing training strategy to narrow the domain gap between synthetic and real‐world datasets. Unlike the previous model [16] that limited the cross-domain data mixing to a single direction, either source-to-target (S2T) or target-to-source (T2S), UDA4Inst combines and balances both strategies to significantly boost instance segmentation performance.
Note that one of the most challenging scenarios in instance segmentation is where multiple instances of the same class overlap. In semantic segmentation, these overlapping instances are treated as a single blob. In contrast, instance segmentation methods need to find clear boundaries in the overlapping region of objects to separate them. We solve this challenge by adopting an instance-wise mixing strategy for large-area instances, ensuring that composite images retain clear object boundaries. For smaller instances, we use a patch-wise mixing method, ensuring that the instances in the mixed images retain the surrounding environmental context. This prevents small, indistinct objects from appearing abruptly or unnaturally in the new domain images. Additionally, to address instance imbalance, UDA4Inst incorporates a straightforward online algorithm that balances rare-class instances during data mixing between domains.
@@
We demonstrate the effectiveness of UDA4Inst on multiple UDA instance segmentation benchmarks. On SYNTHIA [17]Cityscapes, we achieve state-of-the-art results. To the best of our knowledge, this is the first report performing UDA instance segmentation using the synthetic driving scene datasets UrbanSyn [18] and Synscapes [19], achieving 38.6 mAP and 35.9 mAP on Cityscapes [20] and 25.0 mAP and 20.9 mAP on KITTI360 [21]. Compared to the source-only method, our approach demonstrates substantial improvements across these benchmarks as in Fig. 1. Our contributions are as follows:
@@ III-A Overall Approach
Our paper introduces the UDA4Inst framework. Fig. 2 provides an overview of the pipeline. We perform Semantic Category Training by dividing the training data into two semantic groups, (e.g. Human-Cycle and Vehicle). We train an individual model for each group in the Mixing Training Module. Moreover, we apply the bidirectional mixing training, which involves cross-domain cut-and-paste operations performed both instance-wise and patch-wise to train the model. The outputs of these models are subsequently merged and selected by confidence-based assessments to generate high-quality pseudo-labels from the unlabeled target data. Our final step uses only these pseudo-labels to train a single final model and then produce the final predictions in the target domain. This framework aims to generate high-quality pseudo-labels, which are used to train with the target domain data in a supervised manner, ultimately producing the final instance segmentation results. After an overall explanation of our UDA4Inst framework, we explain key proposals in more detail: Semantic Category Training and Bidirectional Mixing Training.
@@ III-C Bidirectional Mixing Training
On each of the two categories, we apply our Bidirectional Mixing Training. The training module pipeline is shown in Fig. 3. For the source domain dataset , comprising images , their instance labels , an instance segmentation model is trained in a supervised manner. The target domain dataset only provides images . We aim to address the domain gap by leveraging , , and to train a model capable of accurately segmenting instances on .
Our teacher-student UDA method is based on online self-training. The training has two stages. In the first stage, the model with parameters is trained on the source domain in a supervised manner for iterations, with synthetic images and their corresponding ground-truth labels , to generate predictions . Our goal is to minimize a supervised segmentation loss
@@ III-C2 Overlapping Strategy
Overlapping instances in the mixing process can create ambiguous labels and unrealistic visual artifacts, potentially confusing the model during training. To address this, we implement a priority masking strategy. Specifically, in S2T mixing, we assign higher priority to the source domain instances. If an instance from the source image overlaps with an instance in the target image, we erase the overlapping region of the target’s instance labels, keeping only the source instance in that region. This means that the source instance’s mask replaces any conflicting target instance masks in the overlapping area. We do an analogous strategy for T2S mixing.
@@ IV-B1 Training
We adopt Mask2Former [15] with Swin Transformer backbone [37] as our base network . The model is pre-trained on COCO [38] dataset. We apply the same data augmentation (flipping, random cropping, etc.) as [15] and set the image crop size to when the target domain is Cityscapes and for KITTI360. For model stabilization, we employ an exponential moving average (EMA) of parameters with a decay rate 0.999. The maximum number of samples N for the rare-class pool is set to 10, and the pseudo-label confidence threshold is 0.9. We set the numbers of iterations for source domain training and mixed data training to be 40k each. During validation, we follow the same checkpoint selection protocol used in [28]. For each stage of training, we set a batch size to be 3, considering optimal memory usage. We use AdamW [39] optimizer with an initial learning rate of 0.0001 and a weight decay of 0.05, as well as the ‘poly’ learning rate schedule with the power of 1. The hyperparameters of the loss , , and are set to 5.0, 5.0, 2.0 respectively. During the supervised training phase using pseudo-labels, we adopt the same training configuration. For bidirectional mixing, we set the pixel area threshold to 1500; instances with an area larger than this threshold are processed instance-wise, while those below are handled patch-wise. For these experiments, UDA4Inst runs on 1 NVIDIA A40 GPU with 48GB of memory.
@@ IV-C1 UDA performance
In Table I, we show how UDA4Inst improves over the vanilla Mask2Former baseline in synth-to-real settings. We also include the supervised Mask2Former (trained on Cityscapes and KITTI360) as upper bounds. UDA4Inst consistently outperforms the source-only model, achieving mAP of 38.6 (UrbanSynCityscapes), 35.9 (SynscapesCityscapes), and a new state-of-the-art 31.3 (SYNTHIACityscapes)—representing +6.6, +7.8, and +6.7 mAP gains over Mask2Former. Improvements are particularly notable for larger objects like Car, Truck, Bus, and Train, while smaller, more deformable classes (Person, Motorcycle, and Bicycle) see modest gains. Fig. 6 further illustrates the advantage of UDA4Inst over the source-only model in UrbanSynCityscapes. UDA4Inst also shows strong performance on KITTI360, reaching 25.0 mAP (UrbanSynKITTI360) and 20.9 mAP (SynscapesKITTI360). Among the source datasets, UrbanSyn consistently performs best on both Cityscapes and KITTI360, likely due to its high photorealism, which narrows the domain gap and boosts adaptation.
@@ IV-D1 Component Ablation
Table III shows how UDA4Inst’s three components—Semantic Category Training (SCT), Bidirectional Mixing Training (BMT), and Rare-class Balancing (RcB)—affect UrbanSynCityscapes. M1 (no UDA) obtains the lowest mAP of 32.0, while M5 (all components) reaches 38.6. Comparing M1 vs. M2 and M3 vs. M4 confirms SCT’s effectiveness, while BMT alone boosts mAP from 32.0 (M1) to 36.1 (M3). With SCT, BMT further improves M2 by +2.2 mAP. Adding RcB (M5) yields substantial gains for rare classes like Train and Motorcycle. Among all components, BMT stands out as the most impactful for adaptation.
@@
w/o mixing | 32.0 |
S2T mixing | 35.4 |
T2S mixing | 35.0 |
bidirectional mixing | 36.1 |
@@ IV-D2 Bidirectional Mixing Training
We examine data mixing strategies without RcB or SCT in Table IV (UrbanSynCityscapes). Both S2T and T2S mixing help, although T2S is slightly weaker due to noise in target pseudo-labels. Combining S2T and T2S achieves 36.1 mAP, a +4.1 gain over the baseline without mixing, balancing target learning with pseudo-label noise mitigation. We also evaluate instance-wise vs. patch-wise mixing in Table V. For the UrbanSyn, Synscapes, and SYNTHIA source datasets, the combined instance‐wise and patch‐wise mixing strategy outperforms either instance‐wise or patch‐wise mixing alone.
@@ IV-D4 Semantic Category Training
Table VII shows that the category approach consistently surpasses the full-class method across all domains. For instance, it yields a +6.3 mAP boost on UrbanSyn and +5.9 on SYNTHIA. Although improvements on Cityscapes are smaller (up to +2.8 mAP), the category model still generalizes better. Similarly, on KITTI360, the category model outperforms the full-class variant. These results confirm that training specialized models for semantically similar classes enhances both source-domain performance and domain adaptation, making semantic grouping a promising strategy for UDA instance segmentation.